OpenStudy (anonymous):

-√147 + √243 simplify

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) The first step would be to find the multiples of 147 and 243. Can you do that for me?

OpenStudy (anonymous):

for 147 its 3 49 and for 243 its 3 and 81

OpenStudy (compassionate):

3 and 49, now 49 can be further factored to 7 * 7, leaving you with: \[-14\sqrt{3}\] Can you do the same for the other one?

OpenStudy (anonymous):

Yea its 3√3

OpenStudy (compassionate):

Okay, so then we have -14 + 3, the radicand stays the same.

OpenStudy (agent0smith):

@Compassionate it should be -7sqrt3 \[\large - \sqrt{7*7*3} = -7\sqrt3\]

OpenStudy (compassionate):

You're right. It would be. So we have -7 + 3

OpenStudy (anonymous):

-4√3

OpenStudy (compassionate):

Correct.

OpenStudy (anonymous):

@agent0smith thats not an answer choice

OpenStudy (agent0smith):

I never said that was the answer.

OpenStudy (compassionate):

Then you got some of your math wrong. Are you sure it's 3[3]? Becuse I took your word for it.

OpenStudy (agent0smith):

It's not \[\large \sqrt{3 * 9*9} = 9 \sqrt3\]

OpenStudy (compassionate):

So then we have -7 + 9

OpenStudy (anonymous):

im sorry i wrot tht down but i must have typed somethin else lol

OpenStudy (compassionate):

Yeah, probably. I suspected it was wrong but I just took your word.

OpenStudy (compassionate):

\[2\sqrt{3}\]

OpenStudy (anonymous):

well thank you lol..and wym that u suspected it was wrong ?

OpenStudy (compassionate):

I suspected you got the 3sqrt{3} thing wrong, but I just went along.

OpenStudy (anonymous):

i wrote down the 9√3 but i just typed it in wrong lol